KIMP Hotel has an expected return on equity of 15% and an after-tax cost of debt of 6%. What debt-equity ratio (i.e., D/E) can produce a WACC of 12%?

0.50

0.67

0.75

0.33
Answer #1

Cost of Equity = ke = 15%

After-tax cost of Debt = kd*(1-t) = 6%

Weight of Equity = We

Weight of Debt = Wd

We + Wd = 1

Wd = 1-We

WACC is calculated using the formula:

WACC = We*ke + Wd*kd*(1-t)

12% = We*15% + (1-We)*6%

12% = We*15% + 6% - We*6%

6% = We*9%

We = 6%/9% = 2/3

We = E/V = 2/3

Wd = D/V = 1/3

Debt-Equity ratio = D/E = Wd/We = (1/3)/(2/3) = 1/2 = 0.5

D/E ratio = 0.5

Answer -> 0.5